OpenStudy (anonymous):

simplify ln (e^3/x^4)

OpenStudy (kinggeorge):

\[\ln \left({e^3 \over x^4}\right) = \ln(e^3) - \ln(x^4) = 3\ln(e) - 4\ln(x)=3-4\ln(x)\]So the final simplified form should be \[3-4\ln(x)\]